OpenStudy (anonymous):

Find the volume of the infinitely long horn that is generated by rotating the region bounded by y = 1 x and y = 0 and lying to the right of x = 1 about the x-axis. where to start just give me a hint.

OpenStudy (anonymous):

hey there im working on another problem. can you follow me.

OpenStudy (anonymous):

well the student isnt responding, ok

OpenStudy (anonymous):

i know i need to write an integration for the disc formed by roation but i don't know how

OpenStudy (anonymous):

waityour problem doesnt make sense

OpenStudy (anonymous):

it says bounded by x =0, y=0, y=1 , x=1 , then to the right of x=1 ?

OpenStudy (anonymous):

Find the volume of the infinitely long horn that is generated by rotating the region bounded by y = 1/x and y = 0 and lying to the right of x = 1 about the x-axis. sorry for the error

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok so integral pi (1/x)^2 from 1 to infinity

OpenStudy (anonymous):

ohk so the infinity part just means that upper limit is infinity :)

OpenStudy (anonymous):

so you get -pi x^-1 from 1 to infinity,

OpenStudy (anonymous):

so -pi / infinity - ( - pi / 1) = 0 + pi = pi

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

make me fan

OpenStudy (anonymous):

already did :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so whats interesting about this, there is no bottom to this horn, so water can travel infinitely, but , the volume is finite?

OpenStudy (anonymous):

i mean if you think of the volume as being submerged in water, you would think, ok , how can this "water bottle horn" have pi units of water